Calculate the standard deviation of these measurements and plot the frequency distribution.


Measurement no.(i) 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15. Diameter d (mm) 1.26 1.26 1.29 1.31 1.28 1.27 1.26 1.25 1.28 1.32 1.21 1.27 1.22 1.29 1.28





1
Expert's answer
2021-05-06T07:49:30-0400

"Mean = \frac{1.26 + 1.26 …. 1.29 + 1.28}{15} = 1.27"

Standard deviations

"σ = \sqrt{\frac{1}{15-1} ((1.26-1.27)^2 + (1.26-1.27)^2 +...+ (1.29-1.27)^2 + (1.28-1.27)^2)} = 0.02927"

Frequency distribution